A newspaper carrier has $3.90 in change. He has three more quarters than dimes but five times as many nickels as quarters. How many coins of each type doe he have?
Answer by stanbon(75887)

A newspaper carrier has $3.90 in change. He has three more quarters than dimes but five times as many nickels as quarters. How many coins of each type does he have?
-----------
5n + 10d + 25q = 390 cents
q = d + 3
n = 5q
------
Substitue for d and n and solve for "q"::
5(5q) + 10(q-3) + 25q = 390
25q + 10q - 30 + 25q = 390
------
60q = 420
---
q = 7 (# of quarters)
d = q-3 = 4 (# of dimes)
n = 5q = 35 (# of nickels)
-------------
Cheers,
Stan H.
------------